MATLAB Primer

Third Edition

Kermit Sigmon Department of Mathematics

University of Florida

Department of Mathematics • University of Florida • Gainesville, FL 32611

sigmon@math.ufl.edu

Copyright \begin{matrix}c \\ ◯ \end{matrix}1989, 1992, 1993 by Kermit Sigmon

On the Third Edition

The Third Edition of the MATLAB Primer is based on version 4.0/4.1 of MATLAB. While this edition reflects an extensive general revision of the Second Edition, most significant is the new information to help one begin to use the major new features of version 4.0/4.1, the sparse matrix and enhanced graphics capabilities.

The plain TEX source and corresponding PostScript file of the latest printing of the MATLAB Primer are always available via anonymous ftp from:

Address: math.ufl.edu  Directory: pub/matlab  Files: primer.tex, primer.ps

You are advised to download anew each term the latest printing of the Primer since minor improvements and corrections may have been made in the interim. If ftp is unavailable to you, the Primer can be obtained via listserv by sending an email message to list-

serv@math.ufl.edu containing the single line send matlab/primer.tex.

Also available at this ftp site are both English (primer35.tex, primer35.ps) and Spanish (primer35sp.tex, primer35sp.ps) versions of the Second Edition of the Primer, which was based on version 3.5 of MATLAB. The Spanish translation is by Celestino Montes, University of Seville, Spain. A Spanish translation of the Third Edition is under development.

Users of the Primer usually appreciate the convenience and durability of a bound copy with a cover, copy center style.

(12-93)

Copyright \begin{matrix}c \\ ◯ \end{matrix}1989, 1992, 1993 by Kermit Sigmon The MATLAB Primer may be distributed as desired subject to the following conditions:

1. It may not be altered in any way, except possibly adding an addendum giving information about the local computer installation or MATLAB toolboxes.

2. It, or any part thereof, may not be used as part of a document distributed for a commercial purpose.

In particular, it may be distributed via a local copy center or bookstore.

Department of Mathematics • University of Florida • Gainesville, FL 32611 sigmon@math.ufl.edu

Introduction

MATLAB is an interactive, matrix-based system for scientific and engineering numeric computation and visualization. You can solve complex numerical problems in a fraction of the time required with a programming language such as Fortran or C. The name MATLAB is derived from MATrix LABoratory.

The purpose of this Primer is to help you begin to use MATLAB. It is not intended to be a substitute for the User’s Guide and Reference Guide for MATLAB. The Primer can best be used hands-on. You are encouraged to work at the computer as you read the Primer and freely experiment with examples. This Primer, along with the on-line help facility, usually suffice for students in a class requiring use of MATLAB.

You should liberally use the on-line help facility for more detailed information. When using MATLAB, the command help functionname will give information about a specific function. For example, the command help eig will give information about the eigenvalue function eig. By itself, the command help will display a list of topics for which on-line help is available; then help topic will list those specific functions under this topic for which help is available. The list of functions in the last section of this Primer also gives most of this information. You can preview some of the features of MATLAB by first entering the command demo and then selecting from the options offered.

The scope and power of MATLAB go far beyond these notes. Eventually you will want to consult the MATLAB User’s Guide and Reference Guide. Copies of the complete documentation are often available for review at locations such as consulting desks, terminal rooms, computing labs, and the reserve desk of the library. Consult your instructor or your local computing center to learn where this documentation is located at your institution.

MATLAB is available for a number of environments: Sun/Apollo/VAXstation/HP workstations, VAX, MicroVAX, Gould, PC and AT compatibles, 80386 and 80486 computers, Apple Macintosh, and several parallel machines. There is a relatively inexpensive Student Edition available from Prentice Hall publishers. The information in these notes applies generally to all of these environments.

MATLAB is licensed by The MathWorks, Inc., 24 Prime Park Way, Natick, MA 01760, (508)653-1415, Fax: (508)653-2997, Email: info@mathworks.com.

Copyright ◯^{c}1989, 1992, 1993 by Kermit Sigmon

Contents

Page

1. Accessing MATLAB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2. Entering matrices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 3. Matrix operations, array operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 4. Statements, expressions, variables; saving a session . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 5. Matrix building functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 6. For, while, if — and relations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7. Scalar functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 8. Vector functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 9. Matrix functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 10. Command line editing and recall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 11. Submatrices and colon notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 12. M-files: script files, function files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 13. Text strings, error messages, input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 14. Managing M-files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 15. Comparing efficiency of algorithms: flops, tic, toc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 16. Output format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 17. Hard copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 18. Graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

planar plots (15), hardcopy (17), 3-D line plots (18) mesh and surface plots (18), Handle Graphics (20)

19. Sparse matrix computations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 20. Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 1. Accessing MATLAB.

On most systems, after logging in one can enter MATLAB with the system command matlab and exit MATLAB with the MATLAB command quit or exit. However, your local installation may permit MATLAB to be accessed from a menu or by clicking an icon.

On systems permitting multiple processes, such as a Unix system or MS Windows, you will find it convenient, for reasons discussed in section 14, to keep both MATLAB and your local editor active. If you are working on a platform which runs processes in multiple windows, you will want to keep MATLAB active in one window and your local editor active in another.

You should consult your instructor or your local computer center for details of the local installation.

2. Entering matrices.

MATLAB works with essentially only one kind of object—a rectangular numerical matrix with possibly complex entries; all variables represent matrices. In some situations, 1-by-1 matrices are interpreted as scalars and matrices with only one row or one column are interpreted as vectors.

Matrices can be introduced into MATLAB in several different ways:

• Entered by an explicit list of elements,

• Generated by built-in statements and functions,

• Created in a diskfile with your local editor,

• Loaded from external data files or applications (see the User’s Guide).

For example, either of the statements

      A = [1 2 3; 4 5 6; 7 8 9]

and

      A = [
      1 2 3
      4 5 6
      7 8 9 ]

creates the obvious 3-by-3 matrix and assigns it to a variable A. Try it. The elements within a row of a matrix may be separated by commas as well as a blank. When listing a number in exponential form (e.g. 2.34e-9), blank spaces must be avoided.

MATLAB allows complex numbers in all its operations and functions. Two convenient ways to enter complex matrices are:

      A = [1 2;3 4] + i*[5 6;7 8]
      A = [1+5i 2+6i;3+7i 4+8i]

When listing complex numbers (e.g. 2+6i) in a matrix, blank spaces must be avoided. Either i or j may be used as the imaginary unit. If, however, you use i and j as variables and overwrite their values, you may generate a new imaginary unit with, say,

ii = sqrt(-1).

Listing entries of a large matrix is best done in an ASCII file with your local editor, where errors can be easily corrected (see sections 12 and 14). The file should consist of a rectangular array of just the numeric matrix entries. If this file is named, say, data.ext (where .ext is any extension), the MATLAB command load data.ext will read this file to the variable data in your MATLAB workspace. This may also be done with a script file (see section 12).

The built-in functions rand, magic, and hilb, for example, provide an easy way to create matrices with which to experiment. The command rand(n) will create an n × n matrix with randomly generated entries distributed uniformly between 0 and 1, while rand(m,n) will create an m × n one. magic(n) will create an integral n × n matrix which is a magic square (rows, columns, and diagonals have common sum); hilb(n) will create the n × n Hilbert matrix, the king of ill-conditioned matrices (m and n denote, of course, positive integers). Matrices can also be generated with a for-loop (see section 6 below).

Individual matrix and vector entries can be referenced with indices inside parentheses in the usual manner. For example, A(2, 3) denotes the entry in the second row, third column of matrix A and x(3) denotes the third coordinate of vector x. Try it. A matrix or a vector will only accept positive integers as indices.

3. Matrix operations, array operations.

The following matrix operations are available in MATLAB:

+  addition
−  subtraction
∗  multiplication
power
′  conjugate transpose
\  left division
/  right division

These matrix operations apply, of course, to scalars (1-by-1 matrices) as well. If the sizes of the matrices are incompatible for the matrix operation, an error message will result, except in the case of scalar-matrix operations (for addition, subtraction, and division as well as for multiplication) in which case each entry of the matrix is operated on by the scalar.

The “matrix division” operations deserve special comment. If A is an invertible square matrix and b is a compatible column, resp. row, vector, then

x = A\b is the solution of A ∗ x = b and, resp., x = b/A is the solution of x ∗ A = b.

In left division, if A is square, then it is factored using Gaussian elimination and these factors are used to solve A ∗ x = b. If A is not square, it is factored using Householder orthogonalization with column pivoting and the factors are used to solve the under- or over- determined system in the least squares sense. Right division is defined in terms of left division by b/A = (A′\b′)′.

Array operations.

The matrix operations of addition and subtraction already operate entry-wise but the other matrix operations given above do not—they are matrix operations. It is important to observe that these other operations, ∗, , \, and /, can be made to operate entry-wise by preceding them by a period. For example, either [1,2,3,4].*[1,2,3,4] or [1,2,3,4]. 2 will yield [1,4,9,16]. Try it. This is particularly useful when using Matlab graphics.

4. Statements, expressions, and variables; saving a session.

MATLAB is an expression language; the expressions you type are interpreted and evaluated. MATLAB statements are usually of the form

variable = expression, or simply expression

Expressions are usually composed from operators, functions, and variable names. Evaluation of the expression produces a matrix, which is then displayed on the screen and assigned to the variable for future use. If the variable name and = sign are omitted, a variable ans (for answer) is automatically created to which the result is assigned.

A statement is normally terminated with the carriage return. However, a statement can be continued to the next line with three or more periods followed by a carriage return. On the other hand, several statements can be placed on a single line if separated by commas or semicolons.

If the last character of a statement is a semicolon, the printing is suppressed, but the assignment is carried out. This is essential in suppressing unwanted printing of intermediate results.

MATLAB is case-sensitive in the names of commands, functions, and variables. For example, solveUT is not the same as solveut.

The command who (or whos) will list the variables currently in the workspace. A variable can be cleared from the workspace with the command clear variablename. The command clear alone will clear all nonpermanent variables.

The permanent variable eps (epsilon) gives the machine unit roundoff—about 10^{−16} on most machines. It is useful in specifying tolerences for convergence of iterative processes.

A runaway display or computation can be stopped on most machines without leaving MATLAB with CTRL-C (CTRL-BREAK on a PC). Saving a session.

When one logs out or exits MATLAB all variables are lost. However, invoking the command save before exiting causes all variables to be written to a non-human-readable diskfile named matlab.mat. When one later reenters MATLAB, the command load will restore the workspace to its former state.

5. Matrix building functions.

Convenient matrix building functions are

eye  identity matrix
zeros  matrix of zeros
ones  matrix of ones
diag  create or extract diagonals
triu  upper triangular part of a matrix
tril  lower triangular part of a matrix
rand  randomly generated matrix
hilb  Hilbert matrix
magic  magic square

                 toeplitz  see help toeplitz

For example, zeros(m,n) produces an m-by-n matrix of zeros and zeros(n) produces an n-by-n one. If A is a matrix, then zeros(size(A)) produces a matrix of zeros having the same size as A.

If x is a vector, diag(x) is the diagonal matrix with x down the diagonal; if A is a square matrix, then diag(A) is a vector consisting of the diagonal of A. What is diag(diag(A))? Try it.

Matrices can be built from blocks. For example, if A is a 3-by-3 matrix, then

      B = [A, zeros(3,2); zeros(2,3), eye(2)]

will build a certain 5-by-5 matrix. Try it.

6. For, while, if — and relations.

In their basic forms, these MATLAB flow control statements operate like those in most computer languages. For.

For example, for a given n, the statement

      x = []; for i = 1:n, x=[x,i 2], end

or

      x = [];
      for i = 1:n
         x = [x,i 2]
      end

will produce a certain n-vector and the statement

      x = []; for i = n:-1:1, x=[x,i 2], end

will produce the same vector in reverse order. Try them. Note that a matrix may be empty (such as x = []).

The statements

for i = 1:m
   for j = 1:n
      H(i, j) = 1/(i+j-1);
   end
end
H

will produce and print to the screen the m-by-n hilbert matrix. The semicolon on the inner statement is essential to suppress printing of unwanted intermediate results while the last H displays the final result.

The for statement permits any matrix to be used instead of 1:n. The variable just consecutively assumes the value of each column of the matrix. For example,

s = 0;
for c = A
   s = s + sum(c);
end

computes the sum of all entries of the matrix A by adding its column sums (Of course, sum(sum(A)) does it more efficiently; see section 8). In fact, since 1:n = [1,2,3,…,n], this column-by-column assigment is what occurs with “if i = 1:n,… ” (see section 11). While.

The general form of a while loop is

while relation

statements

end

The statements will be repeatedly executed as long as the relation remains true. For example, for a given number a, the following will compute and display the smallest nonnegative integer n such that 2^{n} ≥ a:

n = 0;
while 2 n < a
   n = n + 1;
end
n

If.

The general form of a simple if statement is

if relation

statements

end

The statements will be executed only if the relation is true. Multiple branching is also possible, as is illustrated by

if n < 0
   parity = 0;

      elseif rem(n,2) == 0
         parity = 2;
      else
         parity = 1;
      end

In two-way branching the elseif portion would, of course, be omitted. Relations.

The relational operators in MATLAB are

<  less than
>  greater than
<=  less than or equal
>=  greater than or equal
==  equal
∼=  not equal.

Note that “=” is used in an assignment statement while “==” is used in a relation. Relations may be connected or quantified by the logical operators

&  and
|  or
∼  not.

When applied to scalars, a relation is actually the scalar 1 or 0 depending on whether the relation is true or false. Try entering 3 < 5, 3 > 5, 3 == 5, and 3 == 3. When applied to matrices of the same size, a relation is a matrix of 0’s and 1’s giving the value of the relation between corresponding entries. Try a = rand(5), b = triu(a), a == b.

A relation between matrices is interpreted by while and if to be true if each entry of the relation matrix is nonzero. Hence, if you wish to execute statement when matrices A and B are equal you could type

      if A == B

statement

      end

but if you wish to execute statement when A and B are not equal, you would type

      if any(any(A ∼= B))

statement

      end

or, more simply,

      if A == B else

statement

      end

Note that the seemingly obvious

      if A ∼= B, statement, end

will not give what is intended since statement would execute only if each of the corresponding entries of A and B differ. The functions any and all can be creatively used to reduce matrix relations to vectors or scalars. Two any’s are required above since any is a vector operator (see section 8).

7. Scalar functions.

Certain MATLAB functions operate essentially on scalars, but operate element-wise when applied to a matrix. The most common such functions are

sin  asin  exp  abs  round
cos  acos  log (natural log) sqrt  floor
tan  atan  rem (remainder) sign  ceil

8. Vector functions.

Other MATLAB functions operate essentially on a vector (row or column), but act on an m-by-n matrix (m ≥ 2) in a column-by-column fashion to produce a row vector containing the results of their application to each column. Row-by-row action can be obtained by using the transpose; for example, mean(A')'. A few of these functions are

max  sum  median  any
min  prod  mean  all
sort  std

For example, the maximum entry in a matrix A is given by max(max(A)) rather than max(A). Try it.

9. Matrix functions.

Much of MATLAB’s power comes from its matrix functions. The most useful ones are

eig  eigenvalues and eigenvectors
chol  cholesky factorization
svd  singular value decomposition
inv  inverse
lu  LU factorization
qr  QR factorization
hess  hessenberg form
schur  schur decomposition
rref  reduced row echelon form
expm  matrix exponential
sqrtm  matrix square root
poly  characteristic polynomial
det  determinant
size  size
norm  1-norm, 2-norm, F-norm, ∞-norm
cond  condition number in the 2-norm
rank  rank

MATLAB functions may have single or multiple output arguments. For example,

      y = eig(A), or simply eig(A)

produces a column vector containing the eigenvalues of A while

      [U,D] = eig(A)

produces a matrix U whose columns are the eigenvectors of A and a diagonal matrix D with the eigenvalues of A on its diagonal. Try it.

10. Command line editing and recall.

The command line in MATLAB can be easily edited. The cursor can be positioned with the left/right arrows and the Backspace (or Delete) key used to delete the character to the left of the cursor. Other editing features are also available. On a PC try the Home, End, and Delete keys; on a Unix system or a PC the Emacs commands Ctl-a, Ctl-e, Ctl-d, and Ctl-k work; on other systems see help cedit or type cedit.

A convenient feature is use of the up/down arrows to scroll through the stack of previous commands. One can, therefore, recall a previous command line, edit it, and execute the revised command line. For small routines, this is much more convenient that using an M-file which requires moving between MATLAB and the editor (see sections 12 and 14). For example, flopcounts (see section 15) for computing the inverse of matrices of various sizes could be compared by repeatedly recalling, editing, and executing

      a = rand(8); flops(0), inv(a); flops

If one wanted to compare plots of the functions y = sin mx and y = sin nx on the interval [0, 2π] for various m and n, one might do the same for the command line:

      m=2; n=3; x=0:.01:2*pi; y=sin(m*x); z=cos(n*x); plot(x,y,x,z)

11. Submatrices and colon notation.

Vectors and submatrices are often used in MATLAB to achieve fairly complex data manipulation effects. “Colon notation” (which is used both to generate vectors and reference submatrices) and subscripting by integral vectors are keys to efficient manipulation of these objects. Creative use of these features to vectorize operations permits one to minimize the use of loops (which slows MATLAB) and to make code simple and readable. Special effort should be made to become familiar with them.

The expression 1:5 (met earlier in for statements) is actually the row vector [1 2 3 4 5]. The numbers need not be integers nor the increment one. For example,

      0.2:0.2:1.2
gives [0.2, 0.4, 0.6, 0.8, 1.0, 1.2], and
      5:-1:1 gives [5 4 3 2 1].

The following statements will, for example, generate a table of sines. Try it.

      x = [0.0:0.1:2.0]′;
      y = sin(x);
      [x y]

Note that since sin operates entry-wise, it produces a vector y from the vector x.

The colon notation can be used to access submatrices of a matrix. For example, A(1:4,3) is the column vector consisting of the first four entries of the third column of A.

A colon by itself denotes an entire row or column:

A(:,3) is the third column of A, and A(1:4,:) is the first four rows. Arbitrary integral vectors can be used as subscripts:

A(:,[2 4]) contains as columns, columns 2 and 4 of A. Such subscripting can be used on both sides of an assignment statement:

A(:,[2 4 5]) = B(:,1:3) replaces columns 2,4,5 of A with the first three columns of B. Note that the entire altered matrix A is printed and assigned. Try it.

Columns 2 and 4 of A can be multiplied on the right by the 2-by-2 matrix [1 2;3 4]:

      A(:,[2,4]) = A(:,[2,4])*[1 2;3 4]

Once again, the entire altered matrix is printed and assigned.

If x is an n-vector, what is the effect of the statement x = x(n:-1:1)? Try it. Also

try y = fliplr(x) and y = flipud(x').

To appreciate the usefulness of these features, compare these MATLAB statements with a Pascal, FORTRAN, or C routine to effect the same.

12. M-files.

MATLAB can execute a sequence of statements stored in diskfiles. Such files are called “M-files” because they must have the file type of “.m” as the last part of their filename. Much of your work with MATLAB will be in creating and refining M-files. M-files are usually created using your local editor.

There are two types of M-files: script files and function files. Script files.

A script file consists of a sequence of normal MATLAB statements. If the file has the filename, say, rotate.m, then the MATLAB command rotate will cause the statements in the file to be executed. Variables in a script file are global and will change the value of variables of the same name in the environment of the current MATLAB session.

Script files may be used to enter data into a large matrix; in such a file, entry errors can be easily corrected. If, for example, one enters in a diskfile data.m

      A = [
      1 2 3 4
      5 6 7 8
      ];

then the MATLAB statement data will cause the assignment given in data.m to be carried out. However, it is usually easier to use the MATLAB function load (see section 2).

An M-file can reference other M-files, including referencing itself recursively.

Function files.

Function files provide extensibility to MATLAB. You can create new functions specific to your problem which will then have the same status as other MATLAB functions. Variables in a function file are by default local. A variable can, however, be declared global

(see help global).

We first illustrate with a simple example of a function file.

function a = randint(m,n)
%RANDINT Randomly generated integral matrix.
%    randint(m,n) returns an m-by-n such matrix with entries
%    between 0 and 9.
a = floor(10*rand(m,n));

A more general version of this function is the following:

function a = randint(m,n,a,b)
%RANDINT Randomly generated integral matrix.
%    randint(m,n) returns an m-by-n such matrix with entries
%    between 0 and 9.
%    rand(m,n,a,b) return entries between integers a and b.
if nargin < 3, a = 0; b = 9; end
a = floor((b-a+1)*rand(m,n)) + a;

This should be placed in a diskfile with filename randint.m (corresponding to the function name). The first line declares the function name, input arguments, and output arguments; without this line the file would be a script file. Then a MATLAB statement z = randint(4,5), for example, will cause the numbers 4 and 5 to be passed to the variables m and n in the function file with the output result being passed out to the variable z. Since variables in a function file are local, their names are independent of those in the current MATLAB environment.

Note that use of nargin (“number of input arguments”) permits one to set a default value of an omitted input variable—such as a and b in the example.

A function may also have multiple output arguments. For example:

function [mean, stdev] = stat(x)
% STAT Mean and standard deviation
%    For a vector x, stat(x) returns the mean of x;
%    [mean, stdev] = stat(x) both the mean and standard deviation.
%    For a matrix x, stat(x) acts columnwise.
[m n] = size(x);
if m == 1
   m = n; % handle case of a row vector
end
mean = sum(x)/m;
stdev = sqrt(sum(x. 2)/m - mean. 2);

Once this is placed in a diskfile stat.m, a MATLAB command [xm, xd] = stat(x), for example, will assign the mean and standard deviation of the entries in the vector x to xm and xd, respectively. Single assignments can also be made with a function having multiple output arguments. For example, xm = stat(x) (no brackets needed around xm) will assign the mean of x to xm.

The % symbol indicates that the rest of the line is a comment; MATLAB will ignore the rest of the line. Moreover, the first few contiguous comment lines, which document the M-file, are available to the on-line help facility and will be displayed if, for example, help stat is entered. Such documentation should always be included in a function file.

This function illustrates some of the MATLAB features that can be used to produce efficient code. Note, for example, that x. 2 is the matrix of squares of the entries of x, that sum is a vector function (section 8), that sqrt is a scalar function (section 7), and that the division in sum(x)/m is a matrix-scalar operation. Thus all operations are vectorized and loops avoided.

If you can’t vectorize some computations, you can make your for loops go faster by preallocating any vectors or matrices in which output is stored. For example, by including the second statement below, which uses the function zeros, space for storing E in memory is preallocated. Without this MATLAB must resize E one column larger in each iteration, slowing execution.

      M = magic(6);
      E = zeros(6,50);
      for j = 1:50
         E(:,j) = eig(M i);
      end

Some more advanced features are illustrated by the following function. As noted earlier, some of the input arguments of a function—such as tol in this example, may be made optional through use of nargin (“number of input arguments”). The variable nargout can be similarly used. Note that the fact that a relation is a number (1 when true; 0 when false) is used and that, when while or if evaluates a relation, “nonzero” means “true” and 0 means “false”. Finally, the MATLAB function feval permits one to have as an input variable a string naming another function. (Also see eval.)

      function [b, steps] = bisect(fun, x, tol)
      %BISECT Zero of a function of one variable via the bisection method.
      %    bisect(fun,x) returns a zero of the function.  fun is a string
      %    containing the name of a real-valued MATLAB function of a
      %    single real variable; ordinarily functions are defined in
      %    M-files.  x is a starting guess.  The value returned is near
      %    a point where fun changes sign.  For example,
      %    bisect('sin',3) is pi.  Note the quotes around sin.
      %
      %    An optional third input argument sets a tolerence for the
      %    relative accuracy of the result.  The default is eps.
      %    An optional second output argument gives a matrix containing a
      %    trace of the steps; the rows are of form [c f(c)].

% Initialization
if nargin < 3, tol = eps; end
trace = (nargout == 2);
if x ∼= 0, dx = x/20; else, dx = 1/20; end
a = x - dx; fa = feval(fun,a);
b = x + dx; fb = feval(fun,b);
% Find change of sign.
while (fa > 0) == (fb > 0)
   dx = 2.0*dx;
   a = x - dx; fa = feval(fun,a);
   if (fa > 0) ∼= (fb > 0), break, end
   b = x + dx; fb = feval(fun,b);
end
if trace, steps = [a fa; b fb]; end
% Main loop
while abs(b - a) > 2.0*tol*max(abs(b),1.0)
   c = a + 0.5*(b - a); fc = feval(fun,c);
   if trace, steps = [steps; [c fc]]; end
   if (fb > 0) == (fc > 0)
      b = c; fb = fc;
   else
      a = c; fa = fc;
   end
end

Some of MATLAB’s functions are built-in while others are distributed as M-files. The actual listing of any non-built-in M-file—MATLAB’s or your own—can be viewed with the MATLAB command type functionname. Try entering type eig, type vander, and

type rank.

13. Text strings, error messages, input.

Text strings are entered into MATLAB surrounded by single quotes. For example,

s = 'This is a test'

assigns the given text string to the variable s.

Text strings can be displayed with the function disp. For example:

disp('this message is hereby displayed')

Error messages are best displayed with the function error

error('Sorry, the matrix must be symmetric')

since when placed in an M-File, it aborts execution of the M-file.

In an M-file the user can be prompted to interactively enter input data with the function input. When, for example, the statement

      iter = input('Enter the number of iterations:  ')

is encountered, the prompt message is displayed and execution pauses while the user keys in the input data. Upon pressing the return key, the data is assigned to the variable iter and execution resumes.

14. Managing M-files.

While using MATLAB one frequently wishes to create or edit an M-file with the local editor and then return to MATLAB. One wishes to keep MATLAB active while editing a file since otherwise all variables would be lost upon exiting.

This can be easily done using the !-feature. If, while in MATLAB, you precede it with an !, any system command—such as those for editing, printing, or copying a file—can be executed without exiting MATLAB. If, for example, the system command ed accesses your editor, the MATLAB command

      >> !ed rotate.m

will let you edit the file named rotate.m using your local editor. Upon leaving the editor, you will be returned to MATLAB just where you left it.

However, as noted in section 1, on systems permitting multiple processes, such as one running Unix or MS Windows, it may be preferable to keep both MATLAB and your local editor active, keeping one process suspended while working in the other. If these processes can be run in multiple windows, you will want to keep MATLAB active in one window and your editor active in another.

You should consult your instructor or your local computing center for details of the local installation.

Many debugging tools are available. See help dbtype or the list of functions in the last section.

When in MATLAB, the command pwd will return the name of the present working directory and cd can be used to change the working directory. Either dir or ls will list the contents of the working directory while the command what lists only the M-files in the directory. The MATLAB commands delete and type can be used to delete a diskfile and print an M-file to the screen, respectively. While these commands may duplicate system commands, they avoid the use of an !. You may enjoy entering the command why a few times.

M-files must be in a directory accessible to MATLAB. M-files in the present working directory are always accessible. On most mainframe or workstation network installations, personal M-files which are stored in a subdirectory of one’s home directory named matlab will be accessible to MATLAB from any directory in which one is working. The current list of directories in MATLAB’s search path is obtained by the command path. This command can also be used to add or delete directories from the search path. See

help path.

15. Comparing efficiency of algorithms: flops, tic and toc.

Two measures of the efficiency of an algorithm are the number of floating point operations (flops) performed and the elapsed time.

The MATLAB function flops keeps a running total of the flops performed. The command flops(0) (not flops = 0!) will reset flops to 0. Hence, entering flops(0) immediately before executing an algorithm and flops immediately after gives the flop count for the algorithm. For example, the number of flops required to solve a given linear system via Gaussian elimination can be obtained with:

      flops(0), x = A\b; flops

The elapsed time (in seconds) can be obtained with the stopwatch timers tic and toc; tic starts the timer and toc returns the elapsed time. Hence, the commands

tic, any statement, toc will return the elapsed time for execution of the statement. The elapsed time for solving the linear system above can be obtained, for example, with:

      tic, x = A\b; toc

You may wish to compare this time—and flop count—with that for solving the system using x = inv(A)*b;. Try it.

It should be noted that, on timesharing machines, elapsed time may not be a reliable measure of the efficiency of an algorithm since the rate of execution depends on how busy the computer is at the time.

16. Output format.

While all computations in MATLAB are performed in double precision, the format of the displayed output can be controlled by the following commands.

format short  fixed point with 4 decimal places (the default)
format long  fixed point with 14 decimal places
format short e  scientific notation with 4 decimal places
format long e  scientific notation with 15 decimal places
format rat  approximation by ratio of small integers
format hex  hexadecimal format
format bank  fixed dollars and cents
format +  +, -, blank

Once invoked, the chosen format remains in effect until changed.

The command format compact will suppress most blank lines allowing more information to be placed on the screen or page. The command format loose returns to the non-compact format. These commands are independent of the other format commands.

17. Hardcopy.

Hardcopy is most easily obtained with the diary command. The command

diary filename

causes what appears subsequently on the screen (except graphics) to be written to the named diskfile (if the filename is omitted it will be written to a default file named diary) until one gives the command diary off; the command diary on will cause writing to the file to resume, etc. When finished, you can edit the file as desired and print it out on the local system. The !-feature (see section 14) will permit you to edit and print the file without leaving MATLAB.

18. Graphics.

MATLAB can produce planar plots of curves, 3-D plots of curves, 3-D mesh surface plots, and 3-D faceted surface plots. The primary commands for these facilities are plot, plot3, mesh, and surf, respectively. An introduction to each of these is given below.

To preview some of these capabilities, enter the command demo and select some of the graphics options.

Planar plots.

The plot command creates linear x-y plots; if x and y are vectors of the same length, the command plot(x,y) opens a graphics window and draws an x-y plot of the elements of x versus the elements of y. You can, for example, draw the graph of the sine function over the interval -4 to 4 with the following commands:

      x = -4:.01:4; y = sin(x); plot(x,y)

Try it. The vector x is a partition of the domain with meshsize 0.01 while y is a vector giving the values of sine at the nodes of this partition (recall that sin operates entrywise).

You will usually want to keep the current graphics window (“figure”) exposed—but moved to the side—and the command window active.

One can have several graphics figures, one of which will at any time be the designated “current” figure where graphs from subsequent plotting commands will be placed. If, for example, figure 1 is the current figure, then the command figure(2) (or simply figure) will open a second figure (if necessary) and make it the current figure. The command figure(1) will then expose figure 1 and make it again the current figure. The command gcf will return the number of the current figure.

As a second example, you can draw the graph of y = e^{−x^{2}} over the interval -1.5 to 1.5 as follows:

      x = -1.5:.01:1.5; y = exp(-x. 2); plot(x,y)

Note that one must precede by a period to ensure that it operates entrywise (see section 3).

MATLAB supplies a function fplot to easily and efficiently plot the graph of a function. For example, to plot the graph of the function above, one can first define the function in an M-file called, say, expnormal.m containing

function y = expnormal(x)
y = exp(-x. 2);

Then the command

fplot('expnormal', [-1.5,1.5])

will produce the graph. Try it.

Plots of parametrically defined curves can also be made. Try, for example,

t=0:.001:2*pi; x=cos(3*t); y=sin(2*t); plot(x,y)

The graphs can be given titles, axes labeled, and text placed within the graph with the following commands which take a string as an argument.

title  graph title
xlabel  x-axis label
ylabel  y-axis label
gtext  place text on the graph using the mouse
text  position text at specified coordinates

For example, the command

title('Best Least Squares Fit')

gives a graph a title. The command gtext('The Spot') allows one to interactively place the designated text on the current graph by placing the mouse pointer at the desired position and clicking the mouse. To place text in a graph at designated coordinates, one would use the command text (see help text).

The command grid will place grid lines on the current graph. By default, the axes are auto-scaled. This can be overridden by the command axis.

Some features of axis are:

axis([x_{min},x_{max},y_{min},y_{max}])  set axis scaling to prescribed limits
axis(axis)  freezes scaling for subsequent graphs
axis auto  returns to auto-scaling
v = axis  returns vector v showing current scaling
axis square  same scale on both axes
axis equal  same scale and tic marks on both axes
axis off  turns off axis scaling and tic marks
axis on  turns on axis scaling and tic marks

The axis command should be given after the plot command.

Two ways to make multiple plots on a single graph are illustrated by

x=0:.01:2*pi;y1=sin(x);y2=sin(2*x);y3=sin(4*x);plot(x,y1,x,y2,x,y3)

and by forming a matrix Y containing the functional values as columns

x=0:.01:2*pi; Y=[sin(x)', sin(2*x)', sin(4*x)']; plot(x,Y)

Another way is with hold. The command hold on freezes the current graphics screen so that subsequent plots are superimposed on it. The axes may, however, become rescaled. Entering hold off releases the “hold.”

One can override the default linetypes, pointtypes and colors. For example,

      x=0:.01:2*pi; y1=sin(x); y2=sin(2*x); y3=sin(4*x);
      plot(x,y1,'--',x,y2,':',x,y3,'+')

renders a dashed line and dotted line for the first two graphs while for the third the symbol + is placed at each node. The line- and mark-types are

Linetypes: solid (-), dashed (--). dotted (:), dashdot (-.) Marktypes: point (.), plus (+), star (*), circle (o), x-mark (x)

Colors can be specified for the line- and mark-types.

Colors: yellow (y), magenta (m), cyan (c), red (r)

green (g), blue (b), white (w), black (k)

For example, plot(x,y,'r--') plots a red dashed line.

The command subplot can be used to partition the screen so that several small plots can be placed in one figure. See help subplot.

Other specialized 2-D plotting functions you may wish to explore via help are:

      polar, bar, hist, quiver, compass, feather, rose, stairs, fill

Graphics hardcopy

A hardcopy of the current graphics figure can be most easily obtained with the MAT-LAB command print. Entered by itself, it will send a high-resolution copy of the current graphics figure to the default printer.

The printopt M-file is used to specify the default setting used by the print command. If desired, one can change the defaults by editing this file (see help printopt).

The command print filename saves the current graphics figure to the designated filename in the default file format. If filename has no extension, then an appropriate extension such as .ps, .eps, or .jet is appended. If, for example, PostScript is the default file format, then

      print lissajous

will create a PostScript file lissajous.ps of the current graphics figure which can subsequently be printed using the system print command. If filename already exists, it will be overwritten unless you use the -append option. The command

      print -append lissajous

will append the (hopefully different) current graphics figure to the existing file lissajous.ps. In this way one can save several graphics figures in a single file.

The default settings can, of course, be overwritten. For example,

      print -deps -f3 saddle

will save to an Encapsulated PostScript file saddle.eps the graphics figure 3 — even if it is not the current figure.

3-D line plots.

Completely analogous to plot in two dimensions, the command plot3 produces curves in three dimensional space. If x, y, and z are three vectors of the same size, then the command plot3(x,y,z) will produce a perspective plot of the piecewise linear curve in 3-space passing through the points whose coordinates are the respective elements of x, y, and z. These vectors are usually defined parametrically. For example,

      t=.01:.01:20*pi; x=cos(t); y=sin(t); z=t. 3; plot3(x,y,z)

will produce a helix which is compressed near the x-y plane (a “slinky”). Try it.

Just as for planar plots, a title and axis labels (including zlabel) can be added. The features of axis command described there also hold for 3-D plots; setting the axis scaling to prescribed limits will, of course, now require a 6-vector. 3-D mesh and surface plots.

Three dimensional wire mesh surface plots are drawn with the command mesh. The command mesh(z) creates a three-dimensional perspective plot of the elements of the matrix z. The mesh surface is defined by the z-coordinates of points above a rectangular grid in the x-y plane. Try mesh(eye(10)).

Similarly, three dimensional faceted surface plots are drawn with the command surf.

Try surf(eye(10)).

To draw the graph of a function z = f(x, y) over a rectangle, one first defines vectors xx and yy which give partitions of the sides of the rectangle. With the function meshgrid one then creates a matrix x, each row of which equals xx and whose column length is the length of yy, and similarly a matrix y, each column of which equals yy, as follows:

      [x,y] = meshgrid(xx,yy);

One then computes a matrix z, obtained by evaluating f entrywise over the matrices x and y, to which mesh or surf can be applied.

You can, for example, draw the graph of z = e^{−x^{2}−y^{2}} over the square [−2, 2] × [−2, 2] as follows (try it):

      xx = -2:.2:2;
      yy = xx;
      [x,y] = meshgrid(xx,yy);
      z = exp(-x. 2 - y. 2);
      mesh(z)

One could, of course, replace the first three lines of the preceding with

      [x,y] = meshgrid(-2:.2:2, -2:.2:2);

Try this plot with surf instead of mesh.

As noted above, the features of the axis command described in the section on planar plots also hold for 3-D plots as do the commands for titles, axes labelling and the command

hold.

The color shading of surfaces is set by the shading command. There are three settings for shading: faceted (default), interpolated, and flat. These are set by the commands

shading faceted,  shading interp, or shading flat

Note that on surfaces produced by surf, the settings interpolated and flat remove the superimposed mesh lines. Experiment with various shadings on the surface produced above. The command shading (as well as colormap and view below) should be entered after the surf command.

The color profile of a surface is controlled by the colormap command. Available predefined colormaps include:

hsv (default), hot, cool, jet, pink, copper, flag, gray, bone

The command colormap(cool) will, for example, set a certain color profile for the current figure. Experiment with various colormaps on the surface produced above.

The command view can be used to specify in spherical or cartesian coordinates the viewpoint from which the 3-D object is to be viewed. See help view.

The MATLAB function peaks generates an interesting surface on which to experiment

with shading, colormap, and view.

Plots of parametrically defined surfaces can also be made. The MATLAB functions sphere and cylinder will generate such plots of the named surfaces. (See type sphere and type cylinder.) The following is an example of a similar function which generates a plot of a torus.

function [x,y,z] = torus(r,n,a)
%TORUS Generate a torus
%   torus(r,n,a) generates a plot of a torus with central
%   radius a and lateral radius r.  n controls the number
%   of facets on the surface.  These input variables are optional
%   with defaults r = 0.5, n = 30, a = 1.
%
%   [x,y,z] = torus(r,n,a) generates three (n+1)-by-(n+1)
%   matrices so that surf(x,y,z) will produce the torus.
%
%   See also SPHERE, CYLINDER
if nargin < 3, a = 1; end
if nargin < 2, n = 30; end
if nargin < 1, r = 0.5; end
theta = pi*(0:2:2*n)/n;
phi = 2*pi*(0:2:n)'/n;
xx = (a + r*cos(phi))*cos(theta);
yy = (a + r*cos(phi))*sin(theta);
zz = r*sin(phi)*ones(size(theta));
if nargout == 0
   surf(xx,yy,zz)
   ar = (a + r)/sqrt(2);
   axis([-ar,ar,-ar,ar,-ar,ar])
else

         x = xx; y = yy; z = zz;
      end

Other 3-D plotting functions you may wish to explore via help are:

      meshz, surfc, surfl, contour, pcolor

Handle Graphics.

Beyond those described above, MATLAB’s graphics system provides low level functions which permit one to control virtually all aspects of the graphics environment to produce sophisticated plots. Enter the command set(1) and gca,set(ans) to see some of the properties of figure 1 which one can control. This system is called Handle Graphics, for which one is referred to the MATLAB User’s Guide.

19. Sparse Matrix Computations.

In performing matrix computations, MATLAB normally assumes that a matrix is dense; that is, any entry in a matrix may be nonzero. If, however, a matrix contains sufficiently many zero entries, computation time could be reduced by avoiding arithmetic operations on zero entries and less memory could be required by storing only the nonzero entries of the matrix. This increase in efficiency in time and storage can make feasible the solution of significantly larger problems than would otherwise be possible. MATLAB provides the capability to take advantage of the sparsity of matrices.

Matlab has two storage modes, full and sparse, with full the default. The functions full and sparse convert between the two modes. For a matrix A, full or sparse, nnz(A) returns the number of nonzero elements in A.

A sparse matrix is stored as a linear array of its nonzero elements along with their row and column indices. If a full tridiagonal matrix F is created via, say,

      F = floor(10*rand(6)); F = triu(tril(F,1),-1);

then the statement S = sparse(F) will convert F to sparse mode. Try it. Note that the output lists the nonzero entries in column major order along with their row and column indices. The statement F = full(S) restores S to full storage mode. One can check the storage mode of a matrix A with the command issparse(A).

A sparse matrix is, of course, usually generated directly rather than by applying the function sparse to a full matrix. A sparse banded matrix can be easily created via the function spdiags by specifying diagonals. For example, a familiar sparse tridiagonal matrix is created by

      m = 6; n = 6; e = ones(n,1); d = -2*e;
      T = spdiags([e,d,e],[-1,0,1],m,n)

Try it. The integral vector [-1,0,1] specifies in which diagonals the columns of [e,d,e] should be placed (use full(T) to view). Experiment with other values of m and n and, say, [-3,0,2] instead of [-1,0,1]. See help spdiags for further features of spdiags.

The sparse analogs of eye, zeros, ones, and randn for full matrices are, respectively,

      speye,  sparse,  spones,  sprandn

The latter two take a matrix argument and replace only the nonzero entries with ones and normally distributed random numbers, respectively. randn also permits the sparsity structure to be randomized. The command sparse(m,n) creates a sparse zero matrix.

The versatile function sparse permits creation of a sparse matrix via listing its nonzero entries. Try, for example,

      i = [1 2 3 4 4 4]; j = [1 2 3 1 2 3]; s = [5 6 7 8 9 10];
      S = sparse(i,j,s,4,3), full(S)

In general, if the vector s lists the nonzero entries of S and the integral vectors i and j list their corresponding row and column indices, then

      sparse(i,j,s,m,n)

will create the desired sparse m × n matrix S. As another example try

      n = 6; e = floor(10*rand(n-1,1)); E = sparse(2:n,1:n-1,e,n,n)

The arithmetic operations and most MATLAB functions can be applied independent of storage mode. The storage mode of the result? Operations on full matrices always give full results. Selected other results are (S=sparse, F=full):

      Sparse: S+S, S*S, S.*S, S.*F, S n, S. n, S\S
      Full: S+F, S*F, S\F, F\S
      Sparse: inv(S), chol(S), lu(S), diag(S), max(S), sum(S)

For sparse S, eig(S) is full if S is symmetric but undefined if S is unsymmetric; svd requires a full argument. A matrix built from blocks, such as [A,B;C,D], is sparse if any constituent block is sparse.

You may wish to compare, for the two storage modes, the efficiency of solving a tridiagonal system of equations for, say, n = 20, 50, 500, 1000 by entering, recalling and editing the following two command lines:

      n=20;e=ones(n,1);d=-2*e; T=spdiags([e,d,e],[-1,0,1],n,n); A=full(T);
      b=ones(n,1);s=sparse(b);tic,T\s;sparsetime=toc, tic,A\b;fulltime=toc

20. Reference.

There are many MATLAB features which cannot be included in these introductory notes. Listed below are some of the MATLAB functions and operators available, grouped by subject area^{1}. Use the on-line help facility or consult the Reference Guide for more detailed information on the functions.

There are many functions beyond these. There exist, in particular, several “toolboxes” of functions for specific areas^{2}. Included among such are signal processing, control systems, robust-control, system identification, optimization, splines, chemometrics, μ-analysis and synthesis, state-space identification, neural networks, image processing, symbolic math (Maple kernel), and statistics. These can be explored via the command help.

Managing Commands and Functions

help  help facility
what  list M-files on disk
type  list named M-file
lookfor  keywork search through the help entries
which  locate functions and files
demo  run demonstrations
path  control MATLAB’s search path
cedit  set parameters for command line editing and recall
version  display MATLAB version you are running
whatsnew display toolbox README files
info  info about MATLAB and The MathWorks
why  receive flippant answer

Managing Variables and the Workspace

who  list current variables
whos  list current variables, long form
save  save workspace variables to disk
load  retrieve variables from disk
clear  clear variables and functions from memory
pack  consolidate workspace memory
size  size of matrix
length length of vector
disp  display matrix or text

1 Source: MATLAB Reference Guide, version 4.1 2 The toolboxes, which are optional, may not be installed on your system.

Working with Files and the Operating System

cd  change current working directory
pwd  show current working directory
dir, ls  directory listing
delete  delete file
getenv get environment variable
!  execute operating system command
unix  execute operating system command; return result
diary  save text of MATLAB session

Controlling the Command Window

clc  clear command window
home  send cursor home—to top of screen
format set output format
echo  echo commands inside script commands
more  control paged output in command window

Starting and Quitting from MATLAB

quit  terminate MATLAB
startup  M-file executed when MATLAB is started
matlabrc master startup M-file

Matrix Operators  Array Operators
+  addition  + addition
−  subtraction  − subtraction
∗  multiplication  .∗ multiplication
power  .  power
/  right division  ./ right division
\  left division  .\ left division
’  conjugate transpose
.’  transpose
kron Kronecker tensor product

Relational and Logical Operators

<  less than  & and
<= less than or equal  |  or
>  greater than  ∼ not
>= greater than or equal  xor exclusive or
== equal
∼= not equal

Special Characters

=  assignment statement
[ ]  used to form vectors and matrices; enclose multiple function output variables
( ) arithmetic expression precedence; enclose function input variables
.  decimal point
..  parent directory
...  continue statement to next line
,  separate subscripts, function arguments, statements
;  end rows, suppress printing
%  comments
:  subscripting, vector generation
!  execute operating system command

Special Variables and Constraints

ans  answer when expression not assigned
eps  floating point precision
realmax  largest floating point number
reammin  smallest positive floating point number
pi  π
i, j  imaginary unit
inf  infinity
NaN  Not-a-Number
flops  floating point operation count
nargin  number of function input arguments
nargout  number of function output arguments
computer computer type

Time and Date

date  current date
clock  wall clock
etime  elapsed time function
tic, toc  stopwatch timer functions
cputime elapsed CPU time

Special Matrices

zeros  matrix of zeros
ones  matrix of ones
eye  identity
diag  diagonal
toeplitz  Toeplitz
magic  magic square
compan  companion
linspace  linearly spaced vectors
logspace  logarithmically spaced vectors
meshgrid  array for 3-D plots
rand  uniformly distributed random numbers
randn  normally distributed randon numbers
hilb  Hilbert
invhilb  inverse Hilbert (exact)
vander  Vandermonde
pascal  Pascal
hadamard Hadamard
hankel  Hankel
rosser  symmetric eigenvalue test matrix
wilkinson  Wilkinson’s eigenvalue test matrix
gallery  two small test matrices

Matrix Manipulation

diag  create or extract diagonals
rot90  rotate matrix 90 degrees
fliplr  flip matrix left-to-right
flipud  flip matrix up-to-down
reshape change size
tril  lower triangular part
triu  upper triangular part
.’  transpose
:  convert matrix to single column; A(:)

Logical Functions

exist  check if variables or functions exist
any  true if any element of vector is true
all  true if all elements of vector are true
find  find indices of non-zero elements
isnan  true for NaNs
isinf  true for infinite elements
finite  true for finite elements
isieee  true for IEEE floating point arithmetic
isempty true for empty matrix
issparse true for sparse matrix
isstr  true for text string
strcmp  compare string variables

Control Flow

if  conditionally execute statements
else  used with if
elseif  used with if
end  terminate if, for, while
for  repeat statements for a specific number of times
while  repeat statments while condition is true
break terminate execution of for or while loops
return return to invoking function
error  display message and abort function

Programming

input  prompt for user input
keyboard invoke keyboard as if it were a script file
menu  generate menu of choices for user input
pause  wait for user response
function  define function
eval  execute string with MATLAB expression
feval  evaluate function specified by string
global  define global variables
nargchk  validate number of input arguments

Text and Strings

string  about character strings in MATLAB
abs  convert string to numeric values
blanks  a string of blanks
eval  evaluate string with MATLAB expression
num2str convert number to string
int2str  convert integer to string
str2num convert string to number
isstr  true for string variables
strcmp  compare string variables
upper  convert string to uppercase
lower  convert string to lowercase
hex2num convert hex string to floating point number
hex2dec  convert hex string to decimal integer
dec2hex  convert decimal integer to hex string

Debugging

dbstop  set breakpoint
dbclear  remove breakpoint
dbcont  remove execution
dbdown change local workspace context
dbstack  list who called whom
dbstatus list all breakpoints
dbstep  execute one or more lines
dbtype  list M-file with line numbers
dbup  change local workspace context
dbdown opposite of dbup
dbquit  quit debug mode

Sound Processing Functions

saxis  sound axis scaling
sound  convert vector to sound
auread  Read Sun audio file
auwrite Write Sun audio file
lin2mu  linear to mu-law conversion
mu2lin  mu-law to linear conversion

Elementary Math Functions

abs  absolute value or complex magnitude
angle  phase angle
sqrt  square root
real  real part
imag  imaginary part
conj  complex conjugate
gcd  greatest common divisor
lcm  least common multiple
round round to nearest integer
fix  round toward zero
floor  round toward −∞
ceil  round toward ∞
sign  signum function
rem  remainder
exp  exponential base e
log  natural logarithm
log10  log base 10

Trigonometric Functions

sin, asin, sinh, asinh  sine, arcsine, hyperbolic sine, hyperbolic arcsine
cos, acos, cosh, acosh  cosine, arccosine, hyperbolic cosine, hyperbolic arccosine
tan, atan, tanh, atanh tangent, arctangent, hyperbolic tangent, hyperbolic arctangent
cot, acot, coth, acoth  cotangent, arccotangent, hyperbolic cotan., hyperbolic arccotan.
sec, asec, sech, asech  secant, arcsecant, hyperbolic secant, hyperbolic arcsecant
csc, acsc, csch, acsch  cosecant, arccosecant, hyperbolic cosecant, hyperbolic arccosecant

Special Functions

bessel  bessel function
beta  beta function
gamma gamma function
rat  rational approximation
rats  rational output
erf  error function
erfinv  inverse error function
ellipke  complete elliptic integral
ellipj  Jacobian elliptic integral
expint  exponential integral
log2  dissect floating point numbers
pow2  scale floating point numbers

Matrix Decompositions and Factorizations

inv  inverse
lu  factors from Gaussian elimination
rref  reduced row echelon form
chol  Cholesky factorization
qr  orthogonal-triangular decomposition
nnls  nonnegative least squares
lscov  least squares in presence of know covariance
null  null space
orth  orthogonalization
eig  eigenvalues and eigenvectors
hess  Hessenberg form
schur  Schur decomposition
cdf2rdf  complex diagonal form to real block diagonal form
rsf2csf  real block diagonal form to complex diagonal form
balance diagonal scaling for eigenvalue accuracy
qz  generalized eigenvalues
polyeig  polynomial eigenvalue solver
svd  singular value decomposition
pinv  pseudoinverse

Matrix Conditioning

cond  condition number in 2-norm
rcond  LINPACK reciprocal condition number estimator
condest  Hager/Higham condition number estimator
norm  1-norm,2-norm,F-norm,∞-norm
normest 2-norm estimator
rank  rank

Elementary Matrix Functions

expm  matrix exponential
expm1 M-file implementation of expm
expm2 matrix exponential via Taylor series
expm3 matrix exponential via eigenvalues and eigenvectors
logm  matrix logarithm
sqrtm matrix square root
funm  evaluate general matrix function
poly  characteristic polynomial
det  determinant
trace  trace

Polynomials

poly  construct polynomial with specified roots
roots  polynomial roots—companion matrix method
roots1  polynomial roots—Laguerre’s method
polyval  evaluate polynomial
polyvalm evaluate polynomial with matrix argument
conv  multiply polynomials
deconv  divide polynomials
residue  partial-fraction expansion (residues)
polyfit  fit polynomial to data
polyder  differentiate polynomial

Column-wise Data Analysis

max  largest component
min  smallest component
mean  average or mean value
median  median value
std  standard deviation
sort  sort in ascending order
sum  sum of elements
prod  product of elements
cumsum cumulative sum of elements
cumprod cumulative product of elements
hist  histogram

Signal Processing

abs  complex magnitude
angle  phase angle
conv  convolution and polynomial multiplication
deconv  deconvolution and polynomial division
corrcoef  correlation coefficients
cov  covariance matrix
filter  one-dimensional digital filter
filter2  two-dimensional digital filter
cplxpair  sort numbers into complex pairs
unwrap  remove phase angle jumps across 360^{∘} boundaries
nextpow2 next higher power of 2
fft  radix-2 fast Fourier transform
fft2  two-dimensional FFT
ifft  inverse fast Fourier transform
ifft2  inverse 2-D FFT
fftshift  zero-th lag to center of spectrum

Finite Differences and Data Interpolation

diff  approximate derivatives
gradient  approximate gradient
del2  five point discrete Laplacian
subspace angle between two subspaces
spline  cubic spline interpolation
interp1  1-D data interpolation
interp2  2-D data interpolation
interpft  1-D data interpolation via FFT method
griddata  data gridding

Numerical Integration

quad  adaptive 2-panel Simpson’s Rule
quad8 adaptive 8-panel Newton-Cotes Rule
trapz  trapezoidal method

Differential Equation Solution

ode23  2nd/3rd order Runge-Kutta method
ode23p solve via ode23, displaying plot
ode45  4th/5th order Runge-Kutta-Fehlberg method

Nonlinear Equations and Optimization

fmin  minimize function of one variable
fmins minimize function of several variables
fsolve solution to a system of nonlinear equations

(find zeros of a function of several variables)

fzero  find zero of function of one variable
fplot  plot graph of a function

Two Dimensional Graphs

plot  linear plot
loglog  log-log scale plot
semilogx semilog scale plot
semilogy semilog scale plot
fill  draw filled 2-D polygons
polar  polar coordinate plot
bar  bar graph
stairs  stairstep plot
errorbar error bar plot
hist  histogram plot
rose  angle histogram plot
compass compass plot
feather  feather plot
fplot  plot function

Graph Annotation

title  graph title
xlabel  x-axis label
ylabel  y-axis label
zlabel  z-axis label for 3-D plots
grid  grid lines
text  text annotation
gtext  mouse placement of text
ginput graphical input from mouse

Figure Window/Axis Creation and Control

figure  create figure (graph window)
gcf  get handle to current figure
clf  clear current figure
close  close figure
hold  hold current graph
ishold  return hold status
subplot create axes in tiled positions
axes  create axes in arbitrary positions
gca  get handle to to current axes
axis  control axis scaling and appearance
caxis  control pseudocolor axis scaling

Graph Hardcopy and Storage

print  print graph or save graph to file
printopt configure local printer defaults
orient  set paper orientation

Three Dimensional Graphs

mesh  3-D mesh surface
meshc  combination mesh/contour plot
meshz  3-D mesh with zero plane
surf  3-D shaded surface
surfc  combination surface/contour plot
surfl  3-D shaded surface with lighting
plot3  plot lines and points in 3-D space
fill3  draw filled 3-D polygons in 3-D space
contour  contour plot
contour3 3-D contour plot
clabel  contour plot elevation labels
contourc contour plot computation (used by contour)
pcolor  pseudocolor (checkerboard) plot
quiver  quiver plot
image  display image
waterfall waterfall plot
slice  volumetric visualization plot

3-D Graph Appearance

view  3-D graph viewpoint specification
viewmtx view transformation matrices
hidden  mesh hidden line removal mode
shading  color shading mode
axis  axis scaling and apearance
caxis  pseudocolor axis scaling
specular  specular reflectance
diffuse  diffuse reflectance
surfnorm surface normals
colormap color lookup table (see below)
brighten brighten or darken color map
spinmap spin color map
rgbplot  plot colormap
hsv2rgb  hsv to rgb color map conversion
rgb2hsv  rgb to hsv color map conversion

Color Maps

hsv  hue-saturation-value (default)
jet  variant of hsv
gray  linear gray-scale
hot  black-red-yellow-white
cool  shades of cyan and magenta
bone  gray-scale with tinge of blue
copper linear copper tone
pink  pastel shades of pink
flag  alternating red, white, blue, and black

3-D Objects

sphere  generate sphere
cylinder generate cylinder
peaks  generate demo surface

Movies and Animation

moviein  initialize movie frame memory
getframe get movie frame
movie  play recorded movie frames

Handle Graphics Objects

figure  create figure window
axes  create axes
line  create line
text  create text
patch  create patch
surface  create surface
image  create image
uicontrol create user interface control
uimenu  create user interface menu

Handle Graphics Operations

set  set object properties
get  get object properties
reset  reset object properties
delete  delete object
drawnow flush pending graphics events

Sparse Matrix Functions

spdiags  sparse matrix formed from diagonals
speye  sparse identity matrix
sprandn  sparse random matrix
spones  replace nonzero entries with ones
sprandsym sparse symmetric random matrix
spfun  apply function to nonzero entries
sparse  create sparse matrix; convert full matrix to sparse
full  convert sparse matrix to full matrix
find  find indices of nonzero entries
spconvert  convert from sparse matrix external format
issparse  true if matrix is sparse
nnz  number of nonzero entries
nonzeros  nonzero entries
nzmax  amount of storage allocated for nonzero entries
spalloc  allocate memory for nonzero entries
spy  visualize sparsity structure
gplot  plot graph, as in “graph theory”
colmmd  column minimum degree
colperm  order columns based on nonzero count
dmperm  Dulmage-Mendelsohn decomposition
randperm  random permutation vector
symmmd  symmetric minimum degree
symrcm  reverse Cuthill-McKee ordering
condest  estimate 1-norm condition
normest  estimate 2-norm
sprank  structural rank
spaugment form least squares augmented system
spparms  set parameters for sparse matrix routines
symbfact  symbolic factorization analysis
sparsefun  sparse auxillary functions and parameters